# AceFF: A State-of-the-Art Machine Learning Potential for Small Molecules

### v1

@@ 2.5 Drug-like molecules: Schrödinger ligands
The results are plotted in Figure 3a and 3b for AceFF-1.0 and AceFF-2, respectively, showing the force and energy errors. We have plotted guide lines at 0.05eV/Å and 0.05eV, which indicate the approximate validation L1 force and energy loss that the models converge to. We see that, as expected, AceFF-1.0 performs very badly on charged molecules. AceFF-2 performs much better with the majority of force errors below 0.05eV/A. The energy errors are not below 0.05 eV; however, these molecules are mostly larger (range of 27 to 74 atoms) than any it was trained on.
The force errors for all models are plotted in figure 3c. AIMNet2 has consistent performance between charged and neutral molecules, but is systematically above 0.05ev/Å. OrbMol performs exceptionally well with force errors all below 0.03eV/Å. We note that OrbMol was trained on the OMol25 dataset, which does contain molecules of this size (and may well contain these molecules in the training split). It is interesting to see that the most extreme outliers for AIMNet2 are neutral molecules.
@@ 2.6 Potential energy smoothness
All models exhibit similar behavior in the minima. At small distances, we observe the nuclear repulsion in the DFT curve, with the same feature visible in the OrbMol curve (notably, this model includes a ZBL repulsive term). For the other MLIPs, the nuclear repulsion is modeled entirely by the neural network terms (and the data upon which they were trained). The energy barrier in ANI2x is less sharply defined. In contrast, the AceFF-2 energy matches the DFT energy up to a large energy of 40 eV. At the bond-stretching end of the curve, we observe that ANI2x significantly overestimates the bond strength and OrbMol exhibits an unphysical step change between 5 and 6Å, which is perhaps an artifact of the MLIP cutoffs. AIMNet2 and AceFF-2 have similar profiles to each other and to the DFT reference.
@@ 2.7 Inference speed scaling with molecular dynamics
The time per step of all models as a function of system size is shown in Figure 5. An important point to note is that only AceFF and AIMNet2-CUDA graphs utilize CUDA graphs, which are essential for low latency in small systems pelaez2024torchmd.
To make the plots clearer, we have plotted the speed curves in two regimes. In subplot Figure 5a, we see the small system behavior with N below 300, and in subplot Figure 5b, we see the full range of N up to 3000. There is a clear latency limit for very small N for all models except AceFF and AIMNet2-CUDA graphs — the CUDA graph compatibility explained before is the primary reason they can reach near 1ms latency. OrbMol, ANI, and standard AIMNet2 are all limited to >>10ms latency, while the MACE-MPA-A model has even larger latency limits. Once N is large enough, the models scale as expected: Orb, AceFF-1.0, and MACE are linear; AceFF-2 and AIMNet2 are quadratic (due to their $N^{2}$ Coulomb interaction implementations). ANI2X appears to have no clear speed decrease and is limited by latency for all N we are looking at. This follows the speed scaling reported in TorchANI2.0pickering_torchani_2025, where much larger system sizes are reached.
Figure 5: Inference speed of different models vs number of atoms. a. Zoomed in on the x-axis to highlight speed for a small number of atoms. b. Full range of tested atom sizes. Guidelines have been drawn which indicate where 1, 10, and 100 ns/day speeds are if a simulation was run with a 1fs timestep. Methods marked with an * have been run with CUDA graphs enabled. Aimnet2-CUDA graphs marked with † runs out of memory with more than 1500 atoms. This is due to the $N^{2}$ implementation of the CUDA graph-compatible code version.
We also tested the speed of batched molecular dynamics. To do this, we selected a small molecule (imatinib, 68 atoms) and generated N conformers. We then ran Langevin dynamics, using the integrator from TorchMD doerr2021torchmd, in a batched manner, such that all N conformers were integrated simultaneously. TorchMD is a molecular dynamics engine entirely written in PyTorch, which allows for easy integration and batching of molecular simulations using MLIPs and the Amber potential. This allowed the AceFF MLIP to be called once per timestep for the entire batch. The total time per step is plotted in figure 6a for increasing batch size. We compare against the serial OpenMM simulation speed, i.e., the time to sequentially run a timestep for each conformer individually. We also compute the effective total throughput in ns/day for the batched simulations, comparing against the OpenMM simulation throughput of running a single simulation. These results are plotted in figure 6b. We see that for 100 conformers, the batched simulations are approximately 90% faster. The difference is not as dramatic as may be expected, this is because the OpenMM simulations with AceFF-2 can be run with CUDA graphs, which removes nearly all the latency limits that batching typically overcomes.
@@ 2.8 MLIP/MM molecular dynamics stability and speed
A current use-case for MLIPs is in mixed MLIP/MM simulations where a small region is modeled with the MLIP and the rest of the system with a traditional MM forcefield Rufa2020.07.29.227959, sabanes2024enhancing, sabanes_zariquiey_quantumbind-rbfe_2025. The simplest approach is the mechanical embedding approach, where just the ligand intra-molecular forces are computed with the MLIP and the rest of the system uses the MM forcefield. This accounts for the corrections due to ligand strain.
To test that AceFF-2 is appropriate for this, we ran a simulation of the ejm31 ligand bound to Tyk2, which is one of the systems studied in sabanes_zariquiey_quantumbind-rbfe_2025 and from  ross2023maximal, the complex is shown in figure 7a. We used the OpenMM-ML package eastman2023openmm to run the simulations with HMR of 4au, temperature of 300K, and a Langevin thermostat. We ran the simulations for 100ns with a timestep of 1fs and a timestep of 2fs. The ligand RMSDs are plotted in 7b and the center of mass distance between the ligand and protein pocket are plotted in 7c. We observe that both 2fs and 1fs are stable. The simulation speeds on an RTX4090 are 36.7 ns/day and 72.6 ns/day, respectively. We also tested the speed of the same simulation using AceFF-1.0, ANI2x, and MACE-MPA-0 using the implementations available in OpenMM-ML. The speeds for the tested models are shown in table 2 for simulations run with a 1fs timestep.
@@ 2.9 Batched conformer minimization
For AceFF-2, we found that two molecules failed to minimize, resulting in unphysical “exploded” structures. These are molecules with index 1653 (name: GVF_2UVM_A) and index 1816 (name: JD5_4NFK_F). Molecule 1653 has a net charge of -8 due to four -2 phosphate groups; similarly, molecule 1816 has a net charge of -3 and two phosphate groups. Images of these molecules are shown in figure 8b and  8c. The AceFF dataset contains molecules with net charge up to -2 to 2, so these specific molecules are significantly outside the domain of knowledge of the model, and it is not surprising that it fails to correctly model them. The next largest RMSD is 3.9 angstrom for molecule 1928, while not exploded, the conformation is in an elongated state compared to the reference and an H atom has changed bond topology. This molecule has a net charge of -4, so once again outside the training domain. An image of this molecule is shown in figure 8d.
@@
We also assessed the speed of optimizing multiple conformers. The benefit of MLIPs is that batching calculations is trivial using the existing optimizers available in PyTorch. Helpfully, LBFGS, one of the standard methods for geometry optimization, is available. We recorded the time taken to minimize N conformers of Imatinib, a drug molecule with 68 atoms pubchem_imatinib_2025. We varied N from 1 to 100 and put all conformers into the same batch. For comparison, we timed ASE LBFGS minimization (for the same number of steps, in this case 400) for one conformer, and scaled this linearly by N to get the ASE serial speed baseline. The measured timings are shown in Figure 9. The batched speed is significantly faster for more than 10 conformers. In the supplementary information, we verify that LBFGS from PyTorch correctly minimizes atomic configurations in a comparable way to ASE’s LBFGS.
@@ 3 Conclusion
We evaluated AceFF-2 using a comprehensive set of benchmarks, comparing it against other MLIPs and traditional methods. In the current MLIP landscape, ANI-2x ANI2x represents one extreme, offering the fastest speed but with limited elemental coverage and restriction to neutral species. OrbMol orbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, neumann_orb_2024 represents the other extreme, offering high accuracy and coverage of the full OMol25 dataset, spanning most of the periodic table and diverse charge states. AceFF-2 occupies a strategic position between these extremes on the Pareto front; it creates a balance similar in speed and applicability to AIMNet2 anstine2024aimnet2, yet offers superior accuracy.
The TensorNet2 architecture further extends the expressivity of TensorNet and remains one of the fastest equivariant MLIPs available. The models we have primarily compared against are invariant (OrbMol, ANI2x, AIMNet2), these are design choices that may have be influenced by the typically higher speed of invariant models rhodes2025orbv3atomisticsimulationscale when compared to equivariant ones. The only other equivariant model we test is MACE-MPA-0 which is the slowest. We have not investigated direct force prediction MLIPs (i.e., where forces are output in the forward pass and not by back-propagation) as the inherently non-conservative nature adds significant complexity in using them for stable MD simulations bigi2025darkforcesassessingnonconservative.
@@ 4.2 Models and force fields
We used ANI2x ANI2x from TorchANI gao2020torchani. For all evaluations excluding the OpenMM-based ones, we used the latest TorchANI-2 pickering_torchani_2025 (version 2.6) For the OpenMM-based simulation, we used an older version (2.2.4,) which is compatible with the NNPOps galvelis2023nnp (version 0.6) package. For OrbMol-conservative orbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, neumann_orb_2024 we used version 0.5.5 and all calculations were done via the ASE calculator interface. In the text, we refer to the model as OrbMol-conservative and OrbMol interchangeably. We only ever use the conservative model. For GFN2-XTBbannwarth_gfn2-xtbaccurate_2019 we used version 22.1 from the xtb-python package and used the ASE calculator interface for all evaluations. For g-XTBfroitzheim_g-xtb_2025 we used the preliminary version 1.1.0 https://github.com/grimme-lab/g-xtb. We used GAFFwang2004development, wang2006automatic version 2.2 via the OpenMM force-fields package and did all evaluations with OpenMM eastman2023openmm version 8.4. For MACE-MPA-0 we used mace-torch version 0.3.14 and the MPA-0 foundation model batatia2023foundation. We used ASE hjorth_larsen_atomic_2017 version 3.26. We used OpenMM-ML version 1.2, OpenMM-Torch version 1.5, and OpenMM version 8.4.
@@ 4.3 Torsion scans
For the Sellers et al torsion scan, we performed relaxed torsion scans. We used the GeomeTRICwang2016geometry package to perform geometry optimizations with the constrained torsion angles. We took the CCSD(T)/CBS conformation as the starting point for each optimization. To compare with the coupled cluster reference, we computed the relative energy of the torsion scan compared to the minimum energy along the scan. To report a single number from each torsion scan, we computed the mean absolute error of the scan from the reference scan (MAE).
@@ 4.4 Molecular dynamics
For the molecular dynamics in section 2.7, we used the ASE dynamics method. The water boxes were created with OpenMM’s addSolvent functionality. The MD was run with an artificially small timestep and 100 steps. The timings were recorded every 10 steps and the first values were omitted to ignore the overhead of any PyTorch compilation warmup. The simulations were run on an RTX4090. For the batched molecular dynamics, we used a modified version of the Langevin Integrator from TorchMD doerr2021torchmd. For the MLIP/MM molecular dynamics in section 2.8 we used a customized version of the OpenMM-ML package along with OpenMM-Torch version 1.5 and OpenMM version 8.4 on an RTX4090. The simulations were run for 100ns with snapshots saved every 10000 steps. The LangevinMiddle integrator was used with a frequency of 1/ps, a temperature of 300K, HMR of 4au, and a timestep of either 1fs or 2fs. For ANI-2x we employed NNPOpsgalvelis2023nnp. The AceFF simulation was run using torchmd-net configured with max-num-neighbors=32, and CUDA graphs enabled. For MACE-MPA-0 we used the NNPOps neighbor list.
4.5 Updated Triton neighbor list code
In  pelaez2024torchmd we introduced custom CUDA kernels for neighbor list construction. This faster code adds significant complexity to the software build and deployment process, as different versions need to be compiled and released for various PyTorch, Python, CUDA, and platform versions. To ease the development burden, we re-implemented the neighbor list code in Triton, which offers the advantage of being Python-based code that is compiled at runtime by PyTorch. We retained only the ”brute” and ”cell-list” implementations; the brute-force approach is fastest for small systems, while the cell-list is fastest for large systems. We compared the speed of the original CUDA and new Triton versions of both, focusing on a batch size of 1 with an increasing number of atoms. We found that the Triton version is slightly faster. A current limitation of the Triton code is that it cannot yet be exported to TorchScript, which is the primary way to use MLIPs in OpenMM. Therefore, we also implemented a fallback pure-Python neighbor list implementation. The time per step for an increasing number of atoms is plotted in Figure 10. We see that for fewer than 1,000 atoms, all methods exhibit under 1ms latency; for comparison, a 1-layer TensorNet has a latency of at least 20ms for 1,000 atoms. This is an order of magnitude faster. Effectively, neighbor list construction is not a bottleneck for the small MLIP systems we are interested in. For large systems, the Triton cell-list code offers good performance (2ms latency for 100,000 atoms) and could easily be used in other models targeted at these larger system sizes.
Figure 10: Neighbor list speed comparison. The CUDA kernel neighbor list speeds are plotted in blue while the Triton neighbor list speeds are plotted in red. The ”brute” methods are plotted with crosses and the ”cell-list” methods are plotted with circles. The PyTorch baseline is plotted in black, which uses the pure Python fallback code (utilized when the model needs to be torch-scripted). A guideline is plotted at 0.864ms, which corresponds to a molecular dynamics speed of 100ns/day at 1fs.
@@ 4.6 Software and data availability
TensorNet2 is available under the MIT license at https://github.com/torchmd/torchmd-net. We have provided a set of example IPython notebooks (which can be run on Google Colab) and a Python package containing code to enable using AceFF as an ASE (Atomic Simulation Environment hjorth_larsen_atomic_2017) calculator and in OpenMM-MLeastman2023openmm as an ML potential. The package and example notebooks can be found on the GitHub repository https://github.com/Acellera/aceff_examples. The AceFF models can be downloaded from HuggingFace aceff1.1_huggingface. A tutorial to use the models for ML/ML simulations can be found at https://software.acellera.com/acemd/nnpmm.html.
@@
The dataset for the Wiggle150 benchmark can be found at brew_wiggle150_2025. The dataset for the Sellers et al torsion scan can be found at sellers2017comparison. The structures for the Schrodinger Ligand set can be found at ross2023maximal and https://github.com/schrodinger/public_binding_free_energy_benchmark. The Behara et al torsion scan data can be found at behara2024. The Platinum Diverse dataset can be found at friedrich2017.

### v2

@@ 2.6 Drug-like molecules: Schrödinger ligands
The results are plotted in Figure 4a and 4b for AceFF-1.0 and AceFF-2, respectively, showing the force and energy errors. We have plotted guide lines at 0.05eV/Å and 0.05eV, which indicate the approximate validation L1 force and energy loss that the models converge to. We see that, as expected, AceFF-1.0 performs very badly on charged molecules. AceFF-2 performs much better with the majority of force errors below 0.05eV/A. The energy errors are not below 0.05 eV; however, these molecules are mostly larger (range of 27 to 74 atoms) than any it was trained on.
The force errors for all models are plotted in figure 4c. AIMNet2 has consistent performance between charged and neutral molecules, but is systematically above 0.05ev/Å. UMA-S-1.2 performs exceptionally well with force errors all below 0.02eV/Å, closely followed by OrbMol which has all errors below 0.03eV/Å. We note that both OrbMol and UMA-s-1p2 were trained on the OMol25 dataset, which does contain molecules of this size (and may well contain these molecules or very similar ones in the training split). It is interesting to see that the most extreme outliers for AIMNet2 are neutral molecules.
@@ 2.7 Potential energy smoothness
All models exhibit similar behavior in the minima. At small distances, we observe the nuclear repulsion in the DFT curve, with the same feature visible in the OrbMol curve (notably, this model includes a ZBL repulsive term). For the other MLIPs, the nuclear repulsion is modeled entirely by the neural network terms (and the data upon which they were trained). The energy barrier in ANI2x is less sharply defined. In contrast, the AceFF-2 energy matches the DFT energy up to a large energy of 40 eV. At the bond-stretching end of the curve, we observe that ANI2x significantly overestimates the bond strength, UMA-s-1.2 underestimates it, and OrbMol exhibits an unphysical step change between 5 and 6Å, which is perhaps an artifact of the MLIP cutoffs. AIMNet2 and AceFF-2 have similar profiles to each other and to the DFT reference, these two models are the only ones that include a long range coulomb term which enables atom interactions beyond the typical graph cutoff (5Å to 6Å).
@@ 2.8 Inference speed scaling with molecular dynamics
The steps per second of all models as a function of system size is shown in Figure 6. An important point to note is that only AceFF and AIMNet2-CUDA graphs utilize CUDA graphs, which are essential for low latency in small systems 14.
There is a clear latency limit for less than 300 atoms for all models except AceFF and AIMNet2-CUDA graphs — the CUDA graph compatibility explained before is the primary reason they can reach speeds faster than 100 steps/second. OrbMol, ANI, and standard AIMNet2 are all limited to >>10ms latency, while the MACE-MPA-A, and UMA models have even larger latency limits. Once N is large enough, the models scale as expected: Orb, AceFF-1.0, and MACE are linear; AceFF-2 and AIMNet2 are quadratic (due to their $N^{2}$ Coulomb interaction implementations). ANI2X appears to have no clear speed decrease and is limited by latency for all N we are looking at. This follows the speed scaling reported in TorchANI2.0pickering_torchani_2025, where much larger system sizes are reached.
Figure 6: Inference speed in steps per second for different models vs number of atoms. Methods marked with an * have been run with CUDA graphs enabled. AIMNet2-CUDA graphs marked with † runs out of memory with more than 1500 atoms. This is due to the $N^{2}$ implementation of the CUDA graph-compatible code version.
We also tested the speed of batched molecular dynamics. To do this, we selected a small molecule (imatinib, 68 atoms) and generated N conformers. We then ran Langevin dynamics, using the integrator from TorchMD 6, in a batched manner, such that all N conformers were integrated simultaneously. TorchMD is a molecular dynamics engine entirely written in PyTorch, which allows for easy integration and batching of molecular simulations using MLIPs and the Amber potential. This allowed the AceFF MLIP to be called once per timestep for the entire batch. The total time per step is plotted in figure 7a for increasing batch size. We compare against the serial OpenMM simulation speed, i.e., the time to sequentially run a timestep for each conformer individually. We also compute the effective total throughput in ns/day for the batched simulations, comparing against the OpenMM simulation throughput of running a single simulation. These results are plotted in figure 7b. We see that for 32 conformers there is almost a 10x increase in the total simulation throughput, this is due to the latency limits of running a full forward+backwards pass being overcome.
@@ 2.9 MLIP/MM molecular dynamics stability and speed
A current use-case for MLIPs is in mixed MLIP/MM simulations where a small region is modeled with the MLIP and the rest of the system with a traditional MM forcefield 17, 18, sabanes_zariquiey_quantumbind-rbfe_2025. The simplest approach is the mechanical embedding approach, where just the ligand intra-molecular forces are computed with the MLIP and the rest of the system uses the MM forcefield. This accounts for the corrections due to ligand strain.
To test that AceFF-2 is appropriate for this, we ran a simulation of the ejm31 ligand bound to Tyk2, which is one of the systems studied in sabanes_zariquiey_quantumbind-rbfe_2025 and from  16, the complex is shown in figure 8a. We used the OpenMM-ML package 8 to run the simulations with HMR of 4au, temperature of 300K, and a Langevin thermostat. We ran the simulations for 100ns with a timestep of 1fs and a timestep of 2fs. The ligand RMSDs are plotted in 8b and the center of mass distance between the ligand and protein pocket are plotted in 8c. We observe that both 2fs and 1fs are stable. The simulation speeds on an RTX4090 are 36.7 ns/day and 72.6 ns/day, respectively. We also tested the speed of the same simulation using AceFF-1.0, ANI2x, and MACE-MPA-0 using the implementations available in OpenMM-ML. The speeds for the tested models are shown in table 2 for simulations run with a 1fs timestep.
@@ 2.10 Batched conformer minimization
For AceFF-2, we found that two molecules failed to minimize, resulting in unphysical “exploded” structures. These are molecules with index 1653 (name: GVF_2UVM_A) and index 1816 (name: JD5_4NFK_F). Molecule 1653 has a net charge of -8 due to four -2 phosphate groups; similarly, molecule 1816 has a net charge of -3 and two phosphate groups. Images of these molecules are shown in figure 9b and  9c. The AceFF dataset contains molecules with net charge up to -2 to 2, so these specific molecules are significantly outside the domain of knowledge of the model, and it is not surprising that it fails to correctly model them. The next largest RMSD is 3.9 angstrom for molecule 1928, while not exploded, the conformation is in an elongated state compared to the reference and an H atom has changed bond topology. This molecule has a net charge of -4, so once again outside the training domain. An image of this molecule is shown in figure 9d.
@@
We also assessed the speed of optimizing multiple conformers. The benefit of MLIPs is that batching calculations is trivial using the existing optimizers available in PyTorch. Helpfully, LBFGS, one of the standard methods for geometry optimization, is available. We recorded the time taken to minimize N conformers of Imatinib, a drug molecule with 68 atoms pubchem_imatinib_2025. We varied N from 1 to 100 and put all conformers into the same batch. For comparison, we timed ASE LBFGS minimization (for the same number of steps, in this case 400) for one conformer, and scaled this linearly by N to get the ASE serial speed baseline. The measured timings are shown in Figure 10. The batched speed is significantly faster for more than 20 conformers. In the supplementary information, we verify that LBFGS from PyTorch correctly minimizes atomic configurations in a comparable way to ASE’s LBFGS. For fewer than 10 conformers, the ASE version is faster because it utilizes torch.compile and CUDA Graphs to minimize execution latency. Conversely, the batched version used with torch.optim does not employ CUDA Graph settings. We have provided an investigation of speed scaling across various torch.compile configurations in the Supplementary Information.
@@ 3 Conclusion
We evaluated AceFF-2 using a comprehensive set of benchmarks, comparing it against other MLIPs and traditional methods. In the current MLIP landscape, ANI-2x 5 represents one extreme, offering the fastest speed but with limited elemental coverage and restriction to neutral species. OrbMol orbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, neumann_orb_2024 represents the other extreme, offering high accuracy and coverage of the full OMol25 dataset, spanning most of the periodic table and diverse charge states. AceFF-2 occupies a strategic position between these extremes on the Pareto front; it creates a balance similar in speed and applicability to AIMNet2 2, yet offers superior accuracy.
The TensorNet2 architecture further extends the expressivity of TensorNet and remains one of the fastest equivariant MLIPs available. The models we have primarily compared against are invariant (OrbMol, ANI2x, AIMNet2), these are design choices that may have be influenced by the typically higher speed of invariant models rhodes2025orbv3atomisticsimulationscale when compared to equivariant ones. The only other equivariant model we test is MACE-MPA-0 which is significantly slower. We have not investigated direct force prediction MLIPs (i.e., where forces are output in the forward pass and not by back-propagation) as the inherently non-conservative nature adds significant complexity in using them for stable MD simulations bigi2025darkforcesassessingnonconservative.
@@ 4.2 Models and force fields
We used ANI2x 5 from TorchANI 10. For all evaluations excluding the OpenMM-based ones, we used the latest TorchANI-2 pickering_torchani_2025 (version 2.6) For the OpenMM-based simulation, we used an older version (2.2.4,) which is compatible with the NNPOps 9 (version 0.6) package. For OrbMol-conservative orbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, neumann_orb_2024 we used version 0.5.5 and all calculations were done via the ASE calculator interface. In the text, we refer to the model as OrbMol-conservative and OrbMol interchangeably. We only ever use the conservative model. For GFN2-XTBbannwarth_gfn2-xtbaccurate_2019 we used version 22.1 from the xtb-python package and used the ASE calculator interface for all evaluations. For g-XTBfroitzheim_g-xtb_2025 we used the preliminary version 1.1.0 https://github.com/grimme-lab/g-xtb. We used GAFF24, 23 version 2.2 via the OpenMM force-fields package and did all evaluations with OpenMM 8 version 8.4. For MACE-MPA-0 we used mace-torch version 0.3.14 and the MPA-0 foundation model batatia2023foundation. For UMA-s-1.2 we used the omol task from fairchem-core version 2.17. We used ASE hjorth_larsen_atomic_2017 version 3.26. We used OpenMM-ML version 1.2, OpenMM-Torch version 1.5, and OpenMM version 8.4.
@@ 4.3 Torsion scans
For the Sellers et al torsion scan, we performed relaxed torsion scans as done in smith_approaching_2019. We used the GeomeTRIC25 package to perform geometry optimizations with the constrained torsion angles. We took the CCSD(T)/CBS conformation as the starting point for each optimization. To compare with the coupled cluster reference, we computed the relative energy of the torsion scan compared to the minimum energy along the scan. To report a single number from each torsion scan, we computed the mean absolute error of the scan from the reference CCSD(T)/CBS scan (MAE).
@@ 4.4 Molecular dynamics
For the molecular dynamics in section 2.8, we used the ASE dynamics method. The water boxes were created with OpenMM’s addSolvent functionality. The MD was run with an artificially small timestep and 100 steps. The timings were recorded every 10 steps and the first values were omitted to ignore the overhead of any PyTorch compilation warmup. The simulations were run on an RTX4090. For the batched molecular dynamics, we used a modified version of the Langevin Integrator from TorchMD 6. For the MLIP/MM molecular dynamics in section 2.9 we used a customized version of the OpenMM-ML package along with OpenMM-Torch version 1.5 and OpenMM version 8.4 on an RTX4090. The simulations were run for 100ns with snapshots saved every 10000 steps. The LangevinMiddle integrator was used with a frequency of 1/ps, a temperature of 300K, HMR of 4au, and a timestep of either 1fs or 2fs. For ANI-2x we employed NNPOps9. The AceFF simulation was run using torchmd-net configured with max-num-neighbors=32, and CUDA graphs enabled. For MACE-MPA-0 we used the NNPOps neighbor list.
@@ 4.5 Software and data availability
TensorNet2 is available under the MIT license at https://github.com/torchmd/torchmd-net. We have provided a set of example IPython notebooks (which can be run on Google Colab) and a Python package containing code to enable using AceFF as an ASE (Atomic Simulation Environment hjorth_larsen_atomic_2017) calculator and in OpenMM-ML8 as an ML potential. The package and example notebooks can be found on the GitHub repository https://github.com/Acellera/aceff_examples. The AceFF models can be downloaded from HuggingFace aceff1.1_huggingface. A tutorial to use the models for ML/ML simulations can be found at https://software.acellera.com/acemd/nnpmm.html.
@@
The dataset for the Wiggle150 benchmark can be found at brew_wiggle150_2025. The dataset for the Sellers et al torsion scan can be found at 19. The structures for the Schrodinger Ligand set can be found at 16 and https://github.com/schrodinger/public_binding_free_energy_benchmark. The Behara et al torsion scan data can be found at behara2024. The Platinum Diverse dataset can be found at friedrich2017.